OpenStudy (anonymous):

Solve the following system of equations. x - 2y = 5 2x - 4y = 10

OpenStudy (johnweldon1993):

x - 2y = 5 2x - 4y = 10 solve 1 equation for either variable x - 2y = 5 x = 5 + 2y so now we can substitute that into the second equation 2(5 + 2y) - 4y = 10 (see how only "y" is left? that's what we're solving for) 10 + 4y - 4y = 10 10 = 10 y = 0 if this is true...sub it back into the first equation to solve for x x - 2(0) = 5 x - 0 = 5 x = 5 now check both in the second equation 2x - 4y = 10 2(5) - 4(0) = 10 10 - 0 = 10 10 = 10? yes so y = 0 x = 5